OpenStudy (anonymous):

Simplify. 2 times square root 25/16 plz help

OpenStudy (anonymous):

2√(25/16) = 2(5/4) = 10/4 = 5/2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!